# WorldGUI: An Interactive Benchmark for Desktop GUI Automation from Any Starting Point

### v3

@@ 2.1 Task Formulation
Action Space. Our action space includes all raw mouse and keyboard actions, such as left-click, right-click, double-click, drag, keystrokes, and key combinations for shortcuts, among others. Mouse-related actions also specify the target position in the pixel space of the observed screenshot. To ensure a universal and comprehensive representation of actions, we adopted the widely used Python library, PyAutoGUI [^1] https://pyautogui.readthedocs.io, for controlling mouse and keyboard inputs. Each action is represented using the syntax action_type(arguments) as in Table  2.
2.2 Data Collection
2.2.1 Data Source
2.2.2 Pipeline of Data Construction
We engaged four annotators and developed the necessary scripts to structure and format the data. Additionally, to facilitate ground truth (GT) plan generation and pre-action generation, we implemented simple agent-based methods to collect the relevant data. The overall data construction pipeline comprises six steps, as detailed below.
Project File Preparation. Following the AssistGUI assistgui , we create the project file for each task to ensure reproducibility without relying on resource-intensive virtual machines OSWorld or Docker environments windowsagentarena . This approach guarantees that the testing process begins from a consistent state. When combined with pre-actions, it enables augmentation of the same task with various initial states.
GT Plan Generation. We write the script to accept user query $q$ and instructional video $v$ as input and generate the raw plans by agent (powered by GPT-4o). Since the raw plans are not flawless, annotators are asked to watch the videos and manually execute the tasks following the raw plans. During this process, annotators edit the plans to correct any inaccurate steps or descriptions, ultimately producing the finalized GT plans.
[…]
2.3 Evaluation
WorldGUI employs an execution-oriented evaluation approach followed by AssistGUI assistgui and WindowsAgentArena windowsagentarena by utilizing post-processing scripts to assess task completion. Specifically, for tasks like Office work and Web Browsing, we adopt exact matching to compare the differences between the ground-truth (GT) screenshots and the final screenshots. For tasks like File Management, which would produce new folders or change the locations of files, etc. We create the shell script to check the status of files.
Category Applications All Task Task Activities Project File Type
Office PowerPoint 64 Change the content style and layout; Design new effects project.pptx
Office Word 63 Formatting the content style and layout project.docx
Office Excel 70 Table formatting; Data management and processing project.xlsx
Office Adobe Acrobat 66 Automatic add electric signature; Document management project.pdf
Coding VSCode 56 Code editing; Software configuration vscode.exe
Windows Usage Settings 69 Advanced personalized and safety settings; ms-Settings
Windows Usage File Explorer 44 File management: Add, delete, rename, and move files explorer.exe
Web Usage Web Browser 59 Web operation web browser + URL
Web Usage Youtube (Online) 61 Video and account configurations web browser + URL
Media VLC Player 59 Video editing and creation project.mp4
2.4 Data Statistics
WorldGUI compiles GUI tasks from 10 widely used applications on the Windows platform, including productivity software such as PowerPoint, Word, Excel, and VSCode. A total of 111 meta tasks were collected from these applications, with each task being augmented  5 times based on the task’s functionality, resulting in 500 augmented tasks. In total, WorldGUI comprises 611 tasks, and every task has almost 6 variation instances, which is capable of reflecting the real-world interactions of the GUI environment. See the details in Table 3 (more details in the Supplementary Material).
3 WorldGUI-Agent: Thinking before Doing
Figure 3: WorldGUI-Agent. The Planner module receives the user query and an instructional video as input and generates an initial plan. This plan is then refined and executed step by step. Before each step is passed to the Actor module, it undergoes a Step-Check. After the Actor produces an action, the Actor-Critic module iteratively verifies the completion of the action and makes necessary corrections.
In this section, we introduce an universal GUI framework WorldGUI-Agent with a core and essential designing principle: critical thinking, which is vital for designing GUI agents capable of handling dynamic environments that have been overlooked in prior GUI agents  hong2024cogagent ; cheng2024seeclick ; lin2024showui ; zhang2023appagent ; AgentS . The WorldGUI-Agent includes the five fundamental but essential components as in Figure 3 and an Interaction reasoning loop detailed in Algorithm 1. We summarize our critical designs in the following:
@@ 3 WorldGUI-Agent: Thinking before Doing
3.1 State-Aware Planner
The State-Aware Planner processes the instructional video $v$ and the user query $q$ generates an initial plan as shown in the left of Figure 4. We use the speech recognition model Whisper radford2023robust to translate the video $v$ into the subtitle and then send it to the MLLM for task planning. The task plan is hierarchically structured as $p=[p_{1},p_{2},...,p_{N}]$ where $p_{i}$ is a text string describing the $i$-th milestone of the task. Under each $p_{i}$, there is a list of subtasks $[S^{i}_{1},S^{i}_{2},S^{i}_{N}]$, where $S^{i}_{j}$ is the $j$-th subtask in the $i$-th milestone. To ensure the produced plans fit the GUI environment, we propose incorporating an initial screenshot $V_{0}$ to represent the current state. This additional context allows the agent to output plans that align with the actual state. For example, if the instructional video suggests clicking on the “Layout” tab in the Word application, but the current state (as indicated by the screenshot) shows that the “Layout” tab is already selected, there is no need to perform this action again. By utilizing the visual information from the screenshot, the State-Aware Planner can modify the plans accordingly, rather than strictly following the guidance in the instructional video or the existing knowledge from backbone MLLMs. It also avoids the occlusion issue when not seeing the screenshot.
3.2 Planner-Critic
@@
(4) <Reason>: In addition to giving the corrected plans, we force the model to give the reasons. As related works, CoT wei2022chain , OpenAI-o1 openaio1 , and Deepseek-R1 deepseekr1 demonstrate that generating reasoning steps along with the answer would enhance the performance.
3.3 Step-Check
@@
As illustrated in Figure 5, we employ an MLLM to determine whether the current task has been completed or requires modification. We systematically categorize the possible outcomes into four types:
@@
Figure 5: Step-Check. This module first checks the step completion status via an MLLM and then navigates current task processing.
@@
In cases where the screenshot does not provide sufficient visual information for the MLLM to determine the output, the model outputs ##\##Cannot confirm”. When this occurs, we design a Region Search module implemented by an LLM. This module takes the corresponding GUI information extracted by the GUI parser and the task description of the current subtask to identify the relevant region. It then crops the region using the generated bounding box as the center coordinate, with the maximum width and height set to half of the original screenshot dimensions (ensure the region is smaller than the original screenshot). The cropped screenshot is subsequently sent to the Step-Check module to regenerate the decision.
3.4 Actor
@@ 3.5 Actor-Critic
4 Experiments
Method MLLM Office Win. Usage Web Coding Media Overall
@@ 4 Experiments
Plan-Act Gemini-2.0 8.9 3.2 8.3 3.4 28.6 16.2 18.2 2.2 10.0 2.0 6.9
Plan-Act GPT-4o 13.3 10.1 8.3 2.3 23.8 11.1 9.1 2.2 10.0 2.0 8.5
AssistGUI GPT-4o 26.7 16.1 29.2 7.9 33.3 20.2 27.3 11.1 10.0 8.2 16.5
Computer Use Claude-3.5 28.9 19.3 29.2 14.6 71.4 32.3 54.6 22.2 30.0 6.1 23.6
WorldGUI-Agent Gemini-2.0 31.1 17.0 20.8 9.0 38.1 29.3 36.4 11.1 20.0 10.2 19.1
WorldGUI-Agent GPT-4o 42.2 24.3 41.7 11.2 47.6 35.4 45.5 15.6 40.0 12.2 26.0
WorldGUI-Agent Claude-3.5 57.8 32.6 50.0 19.1 76.2 46.5 54.6 26.7 50.0 18.4 36.0
Human 88.9 83.5 100.0 89.9 95.2 80.8 81.8 77.8 90.0 85.7 85.3
Implementation Details. We implement the MLLM in our WorldGUI-Agent by using GPT-4o gpt4o (gpt-4o-2024-08-06) by default. For the computer mouse and keyboard control, we use the Python library PyAutoGUI. Following the AssistGUI assistgui , we use the GUI parser to obtain the position information of elements, e.g., buttons, icons, and text. We use some vision foundation models, such as Google OCR, to extract the text. By default, we use the center coordinates to represent the location of each element. All the testing is under the same screenshot resolution (1920 $\times$ 1080). In all experiments, we set the max trials of the Actor-Critic to 3 for light interaction costs. For the total trials of each task, we set it to $14\times N+$ 1, where $N$ is set empirically.
Evaluation. Given that our WorldGUI includes 611 GUI tasks, we engaged four participants with strong coding and software backgrounds to test all tasks and document their evaluation results. Metric. Following the previous works of OSworld and AssistGUI, we use Success Rate (SR) as the metric.
Baselines. We implement the baseline approach called Plan-Act with different MLLMs as the base model. It focuses on investigating the basic capabilities of task planning and action prediction. Additionally, we compare our WorldGUI-Agent with two strong approaches: AssistGUI assistgui and Computer Use (Claude 3.5) claude3.5 . AssistGUI is a prominent agent framework designed for Desktop GUI Automation, which can plan the task and then execute the task step by step by following the query. We implement it by increasing the MLLM to GPT-4o for better performance. Computer Use (Claude 3.5) is the leading proprietary model specially designed for autonomous computer use. We use the open-source implementation OOTB hu2024dawnguiagentpreliminary as the codebase and then add the subtitle of instructional videos into the input prompt for a fair comparison. We also implement our WorldGUI-Agent with three different MLLMs to illustrate the effectiveness of our proposed universal agent framework.
4.1 Main Results on WorldGUI
Table 4 reports the success rates (SR) of different agents and human experts on our WorldGUI benchmark, broken down by task type (Meta vs. Aug.) across five categories: Office, Win. Usage, Web, Coding, and Media. From these results we draw the following main conclusions.
A large gap remains between agents and humans. The best-performing agent (WorldGUI-Agent with Claude-3.5) achieves an overall SR of only 36.0%, which is less than half of the 85.3% attained by human experts. This stark contrast underscores the difficulty of our tasks and the need for further advances in desktop GUI automation.
Agents generalize poorly to augmented tasks. Across all methods, performance on Augmentation tasks (which introduce interface or context variations) is substantially lower than on their corresponding Meta tasks. For example, Claude-3.5 in the Win. Usage category attains 50.0% on Meta tasks but drops to just 19.1% on Aug. tasks. This highlights the importance of dynamic testing to capture realistic human–computer interaction.
Desktop applications pose a greater challenge than web tasks. Every agent scores higher on Web tasks than on desktop application tasks. WorldGUI-Agent with Claude-3.5, for instance, jumps from 76.2% on Web Meta to only 57.8% on Office Meta, and the gap widens on their Augmentation counterparts. Thus, desktop GUI automation remain a frontier for computer use research.
WorldGUI-Agent consistently outperforms a naive Plan-Act baseline. By incorporating our three critical modules into the planning and execution loop, WorldGUI-Agent substantially improves success rates over the basic Plan-Act approach. Relative to Plan-Act, WorldGUI-Agent raises overall SR by +12.2% with Gemini-2.0, +17.5% with GPT-4o, and +12.4% with Claude-3.5, demonstrating the effectiveness of our design across multiple MLLMs.
Table 5: Success rate (%) of our WorldGUI-Agent (Full Model) with the ablation of different critical modules.
Method Office Win. Usage Web Coding Media Overall
Meta Aug. Meta Aug. Meta Aug. Meta Aug. Meta Aug.
Full Model 42.2 24.3 41.7 11.2 47.6 35.4 45.5 15.6 40.0 12.2 26.0
– w/o Step-Check 31.1 19.3 20.8 9.0 33.3 28.3 45.5 13.3 20.0 8.2 19.8
– w/o Actor-Critic 15.6 7.8 4.2 3.4 28.6 17.2 0.0 8.9 10.0 6.1 9.7
Table 6: Success rate (%) of our WorldGUI-Agent with the ablation of Instructional Video (Inst. Video).
Method PPT Word Excel Acrobat VSCode Overall
Full Model 45.5 36.4 50.0 36.4 45.5 42.9
w/o Inst. Video 45.5 27.3 25.3 18.2 27.3 28.6
4.2 Ablation Study
Impact of different critical modules. Table 5 presents the results of an ablation study on the three core components of WorldGUI-Agent across five application categories (Office, Windows Usage, Web, Coding, Media). The full model achieves an overall success rate (SR) of 26.0%. The effects of removing each component are as follows: Planner-Critic: Eliminating this module reduces overall SR to 18.5% (–7.5%), with substantial drops in Office (42.2% → 31.1%) and Web (47.6% → 38.1%) tasks, indicating its importance for refining initial plans. Step-Check: Without step-wise verification, SR decreases to 19.8% (–6.2%). The relatively smaller decline on Coding and Win. Usage tasks suggest that Step-Check excels at intercepting and correcting multi-step interaction errors. Actor-Critic: Removing the action-level critic causes SR to collapse to 9.7% (–16.3%). Performance on Coding Meta drops to 0.0% and Windows Usage Meta to 4.2%, highlighting the critical role of reward-driven action correction for action-level GUI operations. These results confirm that Planner-Critic, Step-Check, and Actor-Critic each contribute complementary benefits—plan refinement, intermediate validation, and action optimization—that are essential for the robustness and overall effectiveness of WorldGUI-Agent.
Impact of Instructional Video. In Table 6, we study the impact of removing the instructional video by modifying the prompt to include only the user query for generating the initial plan. In the Excel applications, we observe a significant performance decline, as their tasks are complex and difficult, and rely more heavily on additional domain knowledge for successful planning. In contrast, the MLLM performs relatively well on Win. Usage tasks, such as Settings and File Management, are where it has more inherent familiarity. These findings underscore the necessity of instructional videos for complex tasks like visual effect design, mirroring how users learning to build a slide often rely on tutorial videos.
4.3 Results on WindowsAgentArena Benchmark
Table 7 compares WorldGUI-Agent against four leading agents on the WindowsAgentArena benchmark. WorldGUI-Agent achieves a 30.5% overall SR, far surpassing GPT-4V (19.5%), GPT-4o (8.6%), GPT-4o-mini (4.2%), and Phi3-V (3.5%). Its gains are most pronounced in desktop categories: Office tasks (4.7% vs. 0%), Windows System (45.8% vs. 33.3%), and Windows Utilities (33.3% vs. 8.3%). On web browser, it reaches 53.3%, nearly double GPT-4V’s 27.3%, and on coding tasks, it records 33.3% versus 27.3%. In media tasks, WorldGUI-Agent posts 28.6%, closely matching GPT-4V’s 30.3%. These results underscore the necessity of integrated planning critique, step-check verification, and action-level feedback. These results demonstrate that our framework robustly handles both desktop GUI tasks and dynamic web environments, highlighting its versatility for real-world GUI automation.
Table 7: Experimental results on WindowsAgentArena windowsagentarena . The reported results are from the windowsagentarena .
Method Office Web Win. System Coding Media Win. Utils Overall
Phi3-V 0.0 6.9 8.3 0.0 6.2 0.0 3.5
GPT-4o-mini 0.0 14.9 8.3 0.0 0.0 0.0 4.2
GPT-4o 0.0 13.7 29.2 0.0 10.3 0.0 8.6
GPT-4V 0.0 27.3 33.3 27.3 30.3 8.3 19.5
WorldGUI-Agent (Ours) 7.0 53.3 45.8 33.3 28.6 33.3 31.2
Conclusion
In this paper, we take the first step toward comprehensive GUI agent evaluation by introducing WorldGUI. In addition to the standard static testing processes, we incorporate dynamic testing procedures to ensure that WorldGUI effectively captures the complexity and dynamism of real-world GUI environments. Furthermore, we propose a universal agent framework, WorldGUI-Agent, built upon the critical thinking principle. This framework enables the agent to dynamically identify uncommon states and adjust its plans or actions accordingly. Finally, we evaluate WorldGUI-Agent powered by Claude-3.5-Sonnet on WorldGUI and WindowsAgentArena benchmarks, demonstrating the effectiveness across a variety of GUI tasks.
Limitation and Society Impacts
In the current implementation of our agent, WorldGUI-Agent, more external tools have not been integrated into the GUI planning and action prediction processes to prioritize computational efficiency. Incorporating tools such as web search or file search into the agent’s design could be a valuable future direction to improve performance. Additionally, due to the usage of the GUI Parser, which would increase the time costs because of the response speed of experimental desktop computers, it is still a tradeoff between performance and running time in the current GUI domain. We consider that if the base MLLM model is specifically trained with stronger planning and grounding ability, the running time would be sped up. It is noted that our agent framework is capable of working with any MLLM.
WorldGUI takes the first step of pushing the GUI automation into the dynamic testing process, as we found that real-world human-computer interactions are dynamic and unpredictable; existing GUI benchmarks fail to capture such dynamics to closely reflect the interactions. Our WorldGUI-Agent is a straightforward and universal agent framework by considering incorporates three critical modules to adaptively align the plan and actions with exact environment situations, which would be a good baseline for future agent development. For instance, incorporating more tools such as web search or file search into the planning module or action prediction module to realize more challenging tasks.
Appendix A Related Work
A.1 GUI Benchmarks
GUI benchmarks are essential for evaluating the performance and robustness of GUI agents. For web applications, WebShop [29], WebArena [34], and WebVoyager [10] focus on creating the GUI tasks in a web browsing scenario. In OS environments, OSWorld [28] is a comprehensive benchmark, including various operating systems with real applications. Mobile benchmarks such as MobileAgent [23] and AppAgent [31] propose two GUI benchmarks of mobile applications. Windows-related benchmarks like AssistGUI [7] and WindowAgentArena [4] propose a list of real tasks in the Windows platform. However, these online testing GUI benchmarks primarily rely on a static testing process and do not adequately capture the complexity and dynamic nature of GUI environments. As a result, they are insufficient for comprehensively evaluating GUI agents.
A.2 GUI Agents
CogAgent [12] is a vision language model focused on GUI understanding to facilitate GUI navigation, while SeeClick [5] and SeeAct [32] focus on the GUI grounding for enhancing the task performance. MobileAgent [23] and AppAgent [31] are proposed to design the agent on the mobile device. Ferret-UI [30] is another representative work focusing on enhancing the grounding ability in the IOS platform. These agents have shown their ability in GUI understanding (e.g., GUI elements grounding) or action prediction, but still face limitations in handling dynamic and complicated full GUI tasks. Therefore, to enhance GUI automation in dynamic environments, we propose WorldGUI-Agent, which improves adaptability in complex GUI settings and enables agents to effectively handle unpredictable interface changes. The components comparison of our WorldGUI-Agent and other closely related agents is shown in Table 8.
Table 8: Comparison with other closely related agents. Most existing agents solely focus on post-action evaluation but omit the post-planning critique and pre-action validation in handling dynamic GUI environments.
Method Post-Planning Critique Pre-Action Validation Post-Action Evaluation
Mobile-Agent [23]
Mobile-Agent-V2 [23]
AssistGUI [7]
Agent-S [1]
Mobile-Agent-E [24]
WorldGUI-Agent (ours)
A.3 Critical Thinking in Agents
Recent advancements in foundation models and agents, particularly in LLMs such as OpenAI-o1 [19] and Deepseek-R1 [6], have increasingly incorporated thinking processes before providing answers to effectively handle challenging reasoning tasks. The LLM-based agents utilize verify-then-correct process to evaluate and refine intermediate reasoning steps or outputs, ensuring logical coherence and consistency. One notable LLM-based agent framework, Reflexion [22], demonstrates the effectiveness of self-reflection in solving complex tasks. Furthermore, CRITIC [8] integrates external tools into the critique process, leveraging them to improve performance. Noticing that the GUI task is lengthy and complicated, the verify-then-correct process is highly suitable for the GUI scenario. Which is not only aims to enhance the reasoning performance but is also indispensable to designing the key module Actor-Critic [15] to ensure task completion. A closely related work, AssistGUI [7], integrates a critical module only after the Actor module to evaluate action completion. Building upon it, we introduce two additional critical modules: Planner-Critic, applied after the Planner, and Step-Check, applied before the Actor. These two modules lead to a universal and fundamental GUI agent framework WorldGUI-Agent which will provide insights for future GUI agent design.
Appendix B Data
Annotators. In this work, we have four annotators: A, B, C, and D. The team comprises one PhD student, one Master’s student, and two undergraduate students. Prior to annotation, all annotators receive training on using the applications in WorldGUI to ensure high-quality annotations. For the 10 desktop applications, we divide the software into four parts, assigning each part to a different annotator. For the human tests presented in Table 4, the annotators demonstrate tasks on software that they did not annotate. As shown in Table 1, each annotator is responsible for different software during both the annotation and human testing phases to make the soundness of the Human Test results.

### v4

@@ Appendix A Task Formulation
Appendix B WorldGUI-Agent: Thinking before Doing
Figure 6: WorldGUI-Agent. The Planner module receives the user query and an instructional video as input and generates an initial plan. This plan is then refined and executed step by step. Before each step is passed to the Actor module, it undergoes a Step-Check. After the Actor produces an action, the Actor-Critic module iteratively verifies the completion of the action and makes corrections.
In this section, we introduce an universal GUI framework WorldGUI-Agent with a core and essential designing principle: critical thinking, which is vital for designing GUI agents capable of handling dynamic environments that have been overlooked in prior GUI agents  (Hong et al., 2024; Cheng et al., 2024; Lin et al., 2024; Zhang et al., 2023; Agashe et al., 2025a). The WorldGUI-Agent includes the five fundamental but essential components as in Figure 6 and an Interaction reasoning loop detailed in Algorithm 1. We summarize our critical designs in the following:
@@ Appendix B WorldGUI-Agent: Thinking before Doing
B.1 State-Aware Planner
The State-Aware Planner processes the instructional video $v$ and the user query $q$ generates an initial plan as shown in the left of Figure 7. We use the speech recognition model Whisper (Radford et al., 2023) to translate the video $v$ into the subtitle and then send it to the MLLM for task planning. The task plan is hierarchically structured as $p=[p_{1},p_{2},...,p_{N}]$ where $p_{i}$ is a text string describing the $i$-th milestone of the task. Under each $p_{i}$, there is a list of subtasks $S^{i}_{1},S^{i}_{2},S^{i}_{N}]$, where $S^{i}_{j}$ is the $j$-th subtask in the $i$-th milestone. To ensure the produced plans fit the GUI environment, we propose incorporating an initial screenshot $V_{0}$ to represent the current state. This additional context allows the agent to output plans that align with the actual state. For example, if the instructional video suggests clicking on the “Layout” tab in the Word application, but the current state (as indicated by the screenshot) shows that the “Layout” tab is already selected, there is no need to perform this action again. By utilizing the visual information from the screenshot, the State-Aware Planner can modify the plans accordingly, rather than strictly following the guidance in the instructional video or the existing knowledge from backbone MLLMs. It also avoids the occlusion issue when not seeing the screenshot.
B.2 Planner-Critic
@@
(4) <Reason>: In addition to giving the corrected plans, we force the model to give the reasons. As previous studies (e.g., CoT (Wei et al., 2022), Deepseek-R1 (DeepSeek-AI et al., 2025)) demonstrate that generating reasoning steps along with the answer would enhance the performance.
B.3 Step-Check
Figure 8: Step-Check. This module first checks the step completion status via an MLLM and then navigates to the current task processing.
@@ B.3 Step-Check
As illustrated in Figure 8, we employ an MLLM to determine whether the current task has been completed or requires modification. We systematically categorize the possible outcomes into:
@@
In cases where the screenshot does not provide sufficient visual information for the MLLM to determine the output, the model outputs #\#Cannot confirm”. When this occurs, we design a Region Search module implemented by an LLM. This module takes the corresponding GUI information extracted by the GUI parser and the task description of the current subtask to identify the relevant region. It then crops the region using the generated bounding box as the center coordinate, with the maximum width and height set to half of the original screenshot dimensions (ensure the region is smaller than the original screenshot). The cropped screenshot is subsequently sent to the Step-Check module to regenerate the decision.
B.4 Actor
@@ B.5 Actor-Critic
Table 11: Performance comparison between WorldGUI-Agent (with Claude-3.5-Sonnet and Claude-Sonnet-4) and Agent-S2 (Claude-Sonnet-4). Results are reported across five representative applications.
Method PPT VSCode Acrobat VLC File Explorer
@@
Agent-S2 w/ Claude-Sonnet-4 45.5 18.9 45.5 37.8 18.2 14.5 20.0 16.3 60.0 64.7
WorldGUI-Agent w/ Claude-3.5-Sonnet 54.5 39.6 54.5 26.7 63.6 20.0 50.0 18.4 50.0 17.6
WorldGUI-Agent w/ Claude-Sonnet-4 63.6 52.8 54.5 28.9 54.5 30.9 40.0 28.6 70.0 26.5
Table 12: Performance comparison between WorldGUI-Agent (with UI-TARS-1.5) and UI-TARS-1.5.
Method PPT Acrobat
Meta Aug. Meta Aug.
WorldGUI-Agent w/ UI-TARS-1.5 36.6 18.9 36.4 9.1
UI-TARS-1.5 27.3 17.0 9.1 1.8
Appendix C Additional Experiments
As shown in Table 11, Agent-S2 (Agashe et al., 2025b) shows competitive results as compared with our WorldGUI-Agent. We also test on two representative office software to compare the effectiveness of our proposed agentic framework by replacing the base model with UI-TARS-1.5 (Qin et al., 2025) in Table 12. It is noted that to improve the performance of UI-TARS-1.5, we use the GPT-4o to task planning, as we found that UI-TARS struggles with understanding complex desktop software layout and cannot capture the dynamic initial condition changes. We use GPT-4o for better implementation.
Appendix D Data
D.1 Annotators
In this work, we have four annotators: A, B, C, and D. The team comprises one PhD student, one Master’s student, and two undergraduate students. Prior to annotation, all annotators receive training on using the applications in WorldGUI to ensure high-quality annotations. For the 10 desktop applications, we divide the software into four parts, assigning each part to a different annotator. For the human tests presented in Table 4, the annotators demonstrate tasks on software that they did not annotate. As shown in Table 1, each annotator is responsible for different software during both the annotation and human testing phases to make the soundness of the Human Test results.